HETAV PATEL

Grade 11,

At a certain temperature 2 moles of carbon monoxide and 3moles of chlorine were allowed to reach equilibrium according to the reaction CO +Cl2 gives COCl2 in a 5 lit vessel.at equilibrium if one mole of CO is present then equilibrium constant for the reaction is ?

At a certain temperature 2 moles of carbon monoxide and 3moles of chlorine were allowed to reach equilibrium according to the reaction CO +Cl2 gives COCl2 in a 5 lit vessel.at equilibrium if one mole of CO is present then equilibrium constant for the reaction is ?

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
 
CO + Cl2 = CoCl2 2. 3. 0. (At t=0) 2-x. 3-x. x. (At equilibrium)As mentioned in the Q...At equilibrium 1 mole of CO is present therefore...2-x = 1X=1Now... Substituting the value of x we get..At equilibrium...1 mole of CO (1/5). |2 moles of Cl2 (2/5). | (There conc.)1 mole of COCl2 (1/5). |Kc = (1/5) ÷ (2/5)(1/5)= 2.5